OpenStudy (anonymous):

Give Ganeshie8 a medal

OpenStudy (anonymous):

Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. Part 2. Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. Part 3. Explain why the first equation has an extraneous solution and the second does not.

ganeshie8 (ganeshie8):

try this for extraneous solution : \(\large 2\sqrt{x+1}+4 = 2 \) try this for non-extraneous solution : \(\large 2\sqrt{x+1}+2 = 4 \)

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

Wait Just A Quick Question why does the first one have an extraneous solution but the second one doesn't?

ganeshie8 (ganeshie8):

we will get to know that in partB

ganeshie8 (ganeshie8):

solve each equation and see what u get

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

no solution for the first?

OpenStudy (anonymous):

and the second one's solution is 1/4

OpenStudy (anonymous):

right?

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

brb away for lunch forreal this time

ganeshie8 (ganeshie8):

enjoy ur meal :)

OpenStudy (anonymous):

my mom is yelling at me to stop doing math (psh that's a first)

OpenStudy (anonymous):

thx :)

ganeshie8 (ganeshie8):

equation with `extraneous solution` : \(\large 2\sqrt{x+1} + 4 = 2\) \(\large 2\sqrt{x+1} = -2\) \(\large \sqrt{x+1} = -1\) \(\large x+ 1 = 1\) \(\large x = 0\) lets verify if \(x=0\) satisfies the equation : \(\large 2\sqrt{0+1} + 4 = 2\) \(\large 2 + 4 =2\) \(\large 6 = 2\) FALSE. So \(x=0\) does not satisfy this equaion and hence it is an `extraneous solution` for this equation

ganeshie8 (ganeshie8):

similarly you can do the work for solving the second equation

OpenStudy (anonymous):

Oh ok Thank You @ganeshie8